# Adaptive Off-Policy Inference for M-Estimators Under Model Misspecification

### v1

@@ B.1 Proof of Theorem 1
$\dot{s}_{t,\hat{\theta}_{T}}=\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\ddot{m}(X_{t},A_{t},Y_{t})+\pi_{e}(A_{t}=a|X_{t})\sum_{a\in\mathcal{A}}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\ddot{m}(X_{t},a,f_{t}(X_{t},a)).$
It is therefore sufficient to show that
@@
$\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},\right)}\ddot{m}(X_{t},A_{t},Y_{t})\succeq H+o_{p}(1)$
and
@@
$\pi_{e}(A_{t}=a|X_{t})\sum_{a\in\mathcal{A}}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\ddot{m}(X_{t},a,f_{t}(X_{t},a))=o_{p}(1).$
In turn, it is now just required to apply the martingale WLLN because
@@
$\mathbb{E}\left[-\hat{V}_{t}^{-1/2}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},\right)}\ddot{m}(X_{t},A_{t},Y_{t})\right]=\mathbb{E}\left[-\hat{V}_{t}^{-1/2}\mathbb{E}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},\right)}\ddot{m}(X_{t},A_{t},Y_{t})\right]|\mathcal{H}_{t-1}\right]\succeq H,$
by condition 2 of 3. For the second term, the expected value is 0 by the same arguments given above.
@@
By the Cramer-Wold devise, it suffices to show that $\frac{1}{\sqrt{T}}\sum_{t=1}^{T}Z_{t}\overset{d}{\to}N(0,c^{2})$. Consider the martingale difference sequence $Z_{t}:=\frac{1}{\sqrt{T}}c^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}$, for any $c\in\mathbb{R}^{d}$. We now invoke the martingale CLT in Theorem 3.2 of Hall and Heyde (2014). Checking each condition separately:
@@
$\mathbb{E}_{\mathcal{P},p_{t}}[s_{t,\theta^{\star}}|\mathcal{H}_{t-1}]=$ $\mathbb{E}_{\mathcal{P},p_{t}}\left[\mathbb{E}_{\mathcal{P},p_{t}}\left[\frac{\pi_{e}(A_{t}=a|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\dot{m}_{\theta^{\star}}(A_{t},X_{t},Y_{t})|\mathcal{H}_{t-1}\right]|\mathcal{H}_{t-1}\right]$
$+c^{T}\sum_{a=1}^{k}\mathbb{E}_{\mathcal{P},p_{t}}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\right)\dot{m}_{\theta^{\star}}(a,X_{t},f_{t}(a,X_{t})\right].$
@@
$\mathbb{E}_{\mathcal{P},p_{t}}\left[\frac{\pi_{e}(A_{t}=a|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\dot{m}_{\theta^{\star}}(A_{t},X_{t},Y_{t})|\mathcal{H}_{t-1}\right]$ $=\mathbb{E}_{\mathcal{P},\pi_{e}}\left[\dot{m}_{\theta^{\star}}(A_{t},X_{t},Y_{t})|\mathcal{H}_{t-1}\right]=0,$
where the first equality follows because of a change of measure and the last line follows because $\dot{m}_{\theta^{\star}}$ is the minimizer of a score equation.
@@
$\sum_{a=1}^{k}\mathbb{E}_{\mathcal{P},p_{t}}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\dot{m}_{\theta^{\star}}(a,X_{t},f_{t}(a,X_{t})|\mathcal{H}_{t-1}\right]$
$=\sum_{a=1}^{k}\mathbb{E}_{\mathcal{P},p_{t}}\left[\mathbb{E}_{\mathcal{P},p_{t}}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\dot{m}_{\theta^{\star}}(a,X_{t},f_{t}(a,X_{t})|\mathcal{H}_{t-1},X_{t}\right]|\mathcal{H}_{t-1}\right]$
$=\sum_{a=1}^{k}\mathbb{E}_{\mathcal{P},p_{t}}\left[\dot{m}_{\theta^{\star}}(a,X_{t},f_{t}(a,X_{t})\mathbb{E}_{\mathcal{P},p_{t}}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)|\mathcal{H}_{t-1},X_{t}\right]|\mathcal{H}_{t-1}\right].$
However, $\mathbb{E}_{\mathcal{P},p_{t}}\left[1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}|\mathcal{H}_{t-1},X_{t}\right]=0$ since the expectation of the numerartor and denominator are 0. Therefore, both the first and second terms are 0.
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[c^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}\hat{V}_{t}^{-1/2}c\right]$ $=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[c^{T}(\hat{V}_{t}^{-1/2}+V_{t,\theta^{\star}}^{-1/2}-V_{t,\theta^{\star}}^{-1/2})s_{t,\theta^{\star}}\right]$
$=\frac{1}{T}\sum_{t=1}^{T}c^{T}(\hat{V}_{t}^{-1/2}+V_{t,\theta^{\star}}^{-1/2}-V_{t,\theta^{\star}}^{-1/2})\mathbb{E}\left[s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}\right](\hat{V}_{t}^{-1/2}+V_{t,\theta^{\star}}^{-1/2}-V_{t,\theta^{\star}}^{-1/2})c$
$=c^{T}\frac{1}{T}\sum_{t=1}^{T}V_{t,\theta^{\star}}^{-1/2}V_{t,\theta^{\star}}V_{t,\theta^{\star}}^{-1/2}+\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)V_{t,\theta^{\star}}^{1/2}+V_{t,\theta^{\star}}^{1/2}\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)$
$+\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)V_{t,\theta^{\star}}\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)$
@@
where $A_{t}=\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)V_{t,\theta^{\star}}^{1/2}+V_{t,\theta^{\star}}^{1/2}\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)+\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)V_{t,\theta^{\star}}\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)$. By Cesaro summation, it suffices to show that $A_{t}\overset{p}{\to}0$. However, this is true because $V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\overset{p}{\to}0$ by assumption, and $V_{t,\theta^{\star}}^{1/2}$ has bounded eigenvalues.

### v2

@@ B.1 Proof of Theorem 1
$\dot{s}_{t,\hat{\theta}_{T}}=\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\ddot{m}_{\hat{\theta}_{T}}(X_{t},A_{t},Y_{t})+\pi_{e}(A_{t}=a|X_{t})\sum_{a=1}^{K}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\ddot{m}_{\hat{\theta}_{T}}(X_{t},a,f_{t}(X_{t},a)).$
It is sufficient to show the following statements:
  1. 1.
@@
    $\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},\right)}\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})$ is invertible and has bounded eigenvalues with probability tending to 1.
  2. 2.
@@
    $\left\lVert\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},\right)}\left(\ddot{m}_{\hat{\theta}_{T}}(X_{t},A_{t},Y_{t})-\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right)\right\rVert=o_{p}(1)$
  3. 3.
@@
    $\left\lVert\frac{1}{T}\sum_{t=1}^{T}\pi_{e}(A_{t}=a|X_{t})\sum_{a=1}^{K}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\ddot{m}(X_{t},a,f_{t}(X_{t},a))\right\rVert=o_{p}(1).$
Demonstrating (1)
First, note that
$\mathbb{E}\left[\hat{V}_{t}^{-1/2}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},\right)}\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]$ $=\hat{V}_{t}^{-1/2}\mathbb{E}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},\right)}\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]$
$=\hat{V}_{t}^{-1/2}\mathbb{E}_{\mathcal{P},p_{e}}\left[\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right]$
By the weak law of large numbers, we can therefore conclude that
$\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},\right)}\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})$ $=o_{p}(1)+\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\mathbb{E}_{\mathcal{P},p_{e}}\left[\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right].$
Rearranging terms, we can write this as
$o_{p}(1)+\frac{1}{T}\sum_{t=1}^{T}(\hat{V}_{t}^{-1/2}-V_{t,\theta^{\star}}^{1/2})\mathbb{E}_{\mathcal{P},p_{e}}\left[\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right]+\frac{1}{T}\sum_{t=1}^{T}V_{t,\theta^{\star}}^{-1/2}\mathbb{E}_{\mathcal{P},p_{e}}\left[\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right].$
For the second term, we note that:
$\left\lVert\frac{1}{T}\sum_{t=1}^{T}(\hat{V}_{t}^{-1/2}-V_{t,\theta^{\star}}^{1/2})\mathbb{E}_{\mathcal{P},p_{e}}\left[\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right]\right\rVert_{\text{op}}$ $\leq\frac{1}{T}\sum_{t=1}^{T}\left\lVert\hat{V}_{t}^{-1/2}-V_{t,\theta^{\star}}^{1/2}\right\rVert_{\text{op}}\left\lVert\mathbb{E}_{\mathcal{P},p_{e}}\left[\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right]\right\rVert_{\text{op}}$
$=o_{p}(1).$
The last line follows because we know that $\left\lVert\hat{V}_{t}^{-1/2}-V_{t,\theta^{\star}}^{1/2}\right\rVert_{\text{op}}=o_{p}(1)$ and $\left\lVert\mathbb{E}_{\mathcal{P},p_{e}}\left[\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right]\right\rVert_{\text{op}}$ is bounded by 3.
For the final term, we have that $\mathbb{E}_{\mathcal{P},p_{e}}\left[\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right]\succeq H$ for some positive definite matrix $H$ and therefore it has bounded eigenvalues. Similarly, $\left\lVert\frac{1}{T}\sum_{t=1}^{T}V_{t,\theta^{\star}}^{-1/2}\right\rVert_{\text{op}}\leq\delta_{\text{min}}^{-1/2}$ and $\left\lVert\left(\frac{1}{T}\sum_{t=1}^{T}V_{t,\theta^{\star}}^{-1/2}\right)^{-1}\right\rVert_{\text{op}}\geq\delta_{\text{max}}^{-1/2}$. Therefore, the product of both terms has bounded eigenvalues and therefore the third term is invertible with bounded eigenvalues.
Demonstrating (2)
We have by Taylor’s theorem that for some $\tilde{\theta}$ on the line segment connecting $\hat{\theta}_{T}$ and $\tilde{\theta}_{T}$ that,
$\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},\right)}\left(\ddot{m}_{\hat{\theta}_{T}}(X_{t},A_{t},Y_{t})-\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right)$
$=\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},\right)}\dddot{m}_{\tilde{\theta}}(X_{t},A_{t},Y_{t}).$
Taking expectations, we have:
$\left\lVert\mathbb{E}\left[\hat{V}_{t}^{-1/2}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},\right)}\dddot{m}_{\tilde{\theta}}(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]\right\rVert_{1}$ $=\left\lVert\hat{V}_{t}^{-1/2}\mathbb{E}_{\mathcal{P},\pi_{e}}\left[\dddot{m}_{\tilde{\theta}}(X_{t},A_{t},Y_{t})\right]\right\rVert_{1}$
$\leq\left(\left\lVert\hat{V}_{t}^{-1/2}-V_{t,\theta^{\star}}\right\rVert_{1}+\left\lVert V_{t,\theta_{\star}}\right\rVert_{1}\right)\left\lVert\mathbb{E}_{\mathcal{P},\pi_{e}}\left[u_{m}(X_{t},A_{t},Y_{t})\right]\right\rVert_{1}.$
Therefore, by 1 we have that
$\left\lVert\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},\right)}\left(\ddot{m}_{\hat{\theta}_{T}}(X_{t},A_{t},Y_{t})-\ddot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right)\right\rVert$
$\leq o_{p}(1)+\frac{1}{T}\sum_{t=1}^{T}\left(\left\lVert\hat{V}_{t}^{-1/2}-V_{t,\theta^{\star}}^{-1/2}\right\rVert_{1}+\left\lVert V_{t,\theta_{\star}}\right\rVert_{1}\right)\left\lVert\mathbb{E}_{\mathcal{P},\pi_{e}}\left[u_{m}(X_{t},A_{t},Y_{t})\right]\right\rVert_{1}$
$=o_{p}(1)+\frac{1}{T}\sum_{t=1}^{T}\left\lVert V_{t,\theta_{\star}}\right\rVert_{1}\left\lVert\mathbb{E}_{\mathcal{P},\pi_{e}}\left[u_{m}(X_{t},A_{t},Y_{t})\right]\right\rVert_{1}.$
However, the last term is also $o_{p}(1)$ because $V_{t,\theta^{\star}}$ has bounded eigenvalues and $\mathbb{E}_{\mathcal{P},\pi_{e}}\left[u_{m}(X_{t},A_{t},Y_{t})\right]$ is bounded by 7.
Demonstrating (3)
Note that for all $a\in\mathcal{A}$, we have that
$\mathbb{E}\Bigg[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\!\left(A_{t}=a\mid X_{t},\mathcal{H}_{t-1}\right)}\right)\ddot{m}_{\hat{\theta}_{T}}\!\left(X_{t},a,f_{t}(X_{t},a)\right)\,\Bigg|\,\mathcal{H}_{t-1},X_{t}\Bigg]$
$\qquad=\mathbb{E}\Bigg[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\!\left(A_{t}=a\mid X_{t},\mathcal{H}_{t-1}\right)}\right)\,\Bigg|\,\mathcal{H}_{t-1},X_{t}\Bigg]\mathbb{E}\Big[\ddot{m}_{\hat{\theta}_{T}}\!\left(X_{t},a,f_{t}(X_{t},a)\right)\,\Big|\,\mathcal{H}_{t-1},X_{t}\Big]=0.$
Therefore,
$\mathbb{E}\left[\pi_{e}(A_{t}=a|X_{t})\sum_{a=1}^{K}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\ddot{m}_{\hat{\theta}_{T}}(X_{t},a,f_{t}(X_{t},a))|\mathcal{H}_{t-1}\right]=0$
By the martingale WLLN, we then have that
$\frac{1}{T}\sum_{t=1}^{T}\pi_{e}(A_{t}=a|X_{t})\sum_{a=1}^{K}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\ddot{m}(X_{t},a,f_{t}(X_{t},a))=o_{p}(1).$
@@
Consider the term $Z_{t}:=c^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}$, for any $c\in\mathbb{R}^{d}$. By the Cramer-Wold device, it suffices to show that $\frac{1}{\sqrt{T}}\sum_{t=1}^{T}Z_{t}\overset{d}{\to}N(0,c^{2})$. To do this, we demonstrate that $Z_{t}$ is a martingale difference sequence and then apply the martingale CLT of Dvoretzky (1972). In order to apply this fact, we need the following to be true:
  • Conditional Expectation $\mathbb{E}[Z_{t}|\mathcal{H}_{t-1}]=0$ for all $t\in[T]$;
  • Conditional Variance $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Z_{t}^{2}|\mathcal{H}_{t-1}]=\left\lVert c\right\rVert^{2}$;
  • Lindeberg Condition $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Z_{t}^{2}1_{Z_{t}>\epsilon}|\mathcal{H}_{t-1}]\overset{p}{\to}0$.
Checking each condition separately:
@@
$\mathbb{E}[s_{t,\theta^{\star}}|\mathcal{H}_{t-1}]=$ $\mathbb{E}_{\mathcal{P},\pi_{t}}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1},X_{t}\right)}\dot{m}_{\theta^{\star}}(A_{t},X_{t},Y_{t})|\mathcal{H}_{t-1}\right]$
$+\sum_{a=1}^{K}\mathbb{E}_{\mathcal{P},\pi_{t}}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\right)\dot{m}_{\theta^{\star}}(a,X_{t},f_{t}(a,X_{t}))|\mathcal{H}_{t-1}\right].$
@@
$\mathbb{E}_{\mathcal{P},\pi_{t}}\left[\frac{\pi_{e}(A_{t}=a|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\dot{m}_{\theta^{\star}}(A_{t},X_{t},Y_{t})|\mathcal{H}_{t-1}\right]$ $=\mathbb{E}_{\mathcal{P},\pi_{e}}\left[\dot{m}_{\theta^{\star}}(A_{t},X_{t},Y_{t})|\mathcal{H}_{t-1}\right]=0,$
where the first equality follows because of a change of measure and the last line follows because $\theta^{\star}$ is the minimizer of a score equation.
@@
$\sum_{a=1}^{K}\mathbb{E}_{\mathcal{P},\pi_{t}}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\dot{m}_{\theta^{\star}}(a,X_{t},f_{t}(a,X_{t}))|\mathcal{H}_{t-1}\right]$
$=\sum_{a=1}^{K}\mathbb{E}_{\mathcal{P},\pi_{t}}\left[\mathbb{E}_{\mathcal{P},\pi_{t}}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\dot{m}_{\theta^{\star}}(a,X_{t},f_{t}(a,X_{t}))|\mathcal{H}_{t-1},X_{t}\right]|\mathcal{H}_{t-1}\right]$
$=\sum_{a=1}^{K}\mathbb{E}_{\mathcal{P},\pi_{t}}\left[\dot{m}_{\theta^{\star}}(a,X_{t},f_{t}(a,X_{t}))\mathbb{E}_{\mathcal{P},\pi_{t}}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)|\mathcal{H}_{t-1},X_{t}\right]|\mathcal{H}_{t-1}\right].$
However, $\mathbb{E}_{\mathcal{P},\pi_{t}}\left[1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}|\mathcal{H}_{t-1},X_{t}\right]=0$. Therefore, both the first and second terms are 0.
@@
$\frac{1}{T}$ $\sum_{t=1}^{T}\mathbb{E}\left[c^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}\hat{V}_{t}^{-1/2}c|\mathcal{H}_{t-1}\right]$
$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[c^{T}(\hat{V}_{t}^{-1/2}+V_{t,\theta^{\star}}^{-1/2}-V_{t,\theta^{\star}}^{-1/2})s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}(\hat{V}_{t}^{-1/2}+V_{t,\theta^{\star}}^{-1/2}-V_{t,\theta^{\star}}^{-1/2})c|\mathcal{H}_{t-1}\right]$
$=\frac{1}{T}\sum_{t=1}^{T}c^{T}(\hat{V}_{t}^{-1/2}+V_{t,\theta^{\star}}^{-1/2}-V_{t,\theta^{\star}}^{-1/2})\mathbb{E}\left[s_{t,\theta^{\star}}s_{t,\theta^{\star}}|\mathcal{H}_{t-1}^{T}\right](\hat{V}_{t}^{-1/2}+V_{t,\theta^{\star}}^{-1/2}-V_{t,\theta^{\star}}^{-1/2})c$
$=c^{T}\frac{1}{T}\sum_{t=1}^{T}\Big[V_{t,\theta^{\star}}^{-1/2}V_{t,\theta^{\star}}V_{t,\theta^{\star}}^{-1/2}-\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)V_{t,\theta^{\star}}^{1/2}-V_{t,\theta^{\star}}^{1/2}\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)$
$\hskip 60.00009pt+\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)V_{t,\theta^{\star}}\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)\Big]c$
@@
where $A_{t}=\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)V_{t,\theta^{\star}}^{1/2}+V_{t,\theta^{\star}}^{1/2}\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)+\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)V_{t,\theta^{\star}}\left(V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right)$. By Cesaro summation, it suffices to show that $\left\lVert A_{t}\right\rVert_{\text{op}}\overset{p}{\to}0$. However, this is true because $\left\lVert V_{t,\theta^{\star}}^{-1/2}-\hat{V}_{t}^{-1/2}\right\rVert_{\text{op}}\overset{p}{\to}0$ by assumption, and $V_{t,\theta^{\star}}^{1/2}$ has bounded eigenvalues.